# Chains of model structures arising from modules of finite Gorenstein dimension

### v3

@@ Abstract.
Keywords: projectively coresolved Gorenstein flat (PGF) module; module of finite PGF (Gorenstein projective, Gorenstein flat) dimension; (complete, hereditary) cotorsion pair; abelian (exact) model structure; weakly idempotent exact category; finitistic dimension
@@ Introduction
This kind of PGF modules enjoys pleasant properties. It is not clear whether a Gorenstein projective module is Gorenstein flat. But PGF modules provide a common refinement of Gorenstein projective modules and Gorenstein flat modules: by definition they are Gorenstein flat, and they also prove to be Gorenstein projective ([28, Theorem 4.4]). The full subcategory $\mathcal{PGF}$ of $R$-Mod consisting of the PGF modules is a Frobenius category with $\mathcal{P}$, the full subcategory of projective $R$-modules, as the class of projective-injective objects, and the stable triangulated category $\mathcal{PGF}/\mathcal{P}$ is the homotopy category of infinite different abelian or exact model structures, in the sense of D. Quillen [24], M. Hovey [19], and J. Gillespie [14].
What interesting and important are, all the three classes $\mathcal{PGF}$, $\mathcal{GP}$, and $\mathcal{GF}$, are the left parts of cotorsion pairs in $R$-Mod, where $\mathcal{GP}$ (respectively, $\mathcal{GF}$) is the full subcategory of $R$-Mod consisting of the Gorenstein projective (respectively, Gorenstein flat) modules. A. Beligiannis and I. Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra. Šaroch and Št́ovíček [28, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$. By a result of A. Iacob [20, Proposition 9] (see also [22, Proposition 4.11]), if $R$ is a Gorenstein ring (i.e., $R$ is a left and right noetherian ring and the injective dimensions of ${}_{R}R$ and $R_{R}$ are finite), then $\mathcal{PGF}=\mathcal{GP}$. Thus, in this case, the cotorsion pair $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ reads as $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, which has been given by Hovey [19, Theorem 8.3]. By the Hovey correspondence, all these three kinds of cotorsion pairs induce abelian model structures in $R$-Mod.
@@
are complete and hereditary cotorsion pairs in $R$-Mod (see [11, Theorem 7.4.6], and [23, Theorem 3.4(2)], respectively), where $\mathcal{P}_{n}$ (respectively, $\mathcal{F}_{n})$ is the class of $R$-modules of projective (respectively, flat) dimension $\leq n$.
@@
is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra, where $\mathcal{GP}_{n}$ is the class of $R$-modules of Gorenstein projective dimension $\leq n$ (see Theorem 3.4). R. El Maaouy [22, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
@@
Enochs, Jenda [11, Theorem 7.4.6]
@@
Enochs, Jenda [11, Proposition 7.4.3]
@@ 1.1. Modules of finite Gorenstein dimension
Following [28], an $R$-module $M$ is projectively coresolved Gorenstein flat (PGF for short), if it is a syzygy of an exact sequence of projective modules which remains exact after tensoring with any right injective $R$-module. Denote by $\mathcal{PGF}$ the subcategory of PGF modules. By definition $\mathcal{PGF}\subseteq\mathcal{GF}.$ It is proved in [28, Theorem 4.4] that $\mathcal{PGF}\subseteq\mathcal{GP}.$ However, it is not clear whether a Gorenstein projective module is PGF. In fact, $\mathcal{PGF}=\mathcal{GP}$ if and only if $\mathcal{GP}\subseteq\mathcal{GF}$. Denoted by ${\rm PGFd}M$ the PGF dimension of module $M$ (see [8]). Denote by $\mathcal{PGF}^{<\infty}$ the class of modules of finite PGF dimension. For a non-negative integer $n$, let $\mathcal{PGF}_{n}$ be the subcategory of $R$-modules $M$ with ${\rm PGFd}M\leq n$. Thus one has
@@ 1.2. Cotorsion pairs in exact categories
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$, where $\mathcal{S}$ is a set of modules (or, equivalently, cogenerated by a single module). Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([12, Theorem 10]; also [16, Theorem 6.11(b)]: but in [16] “cogenerated” is called “generated”). This result has been generalized to Grothendieck category with enough projective objects in [19, Theorem 2.4].
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is hereditary, if $\mathcal{C}$ is closed under the kernel of deflations, and $\mathcal{F}$ is closed under the cokernel of inflations. For a cotorsion pair $(\mathcal{C},\mathcal{F})$ in an abelian category $\mathcal{A}$ with enough projective objects and injective objects, these two conditions of the heredity are equivalent; and they are equivalent to ${{\rm Ext}}^{2}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$ and $F\in\mathcal{F};$ and also to ${{\rm Ext}}^{i}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$, $F\in\mathcal{F}$, and $i\geq 2$. See [13, Proposition 1.2.10]. This result holds also for a complete cotorsion pair in a weakly idempotent complete exact category. See [29, Lemma 6.17] (also [27]).
@@ 1.3. Model structures
For a model structure $(\operatorname{CoFib}$,  $\operatorname{Fib}$,  $\operatorname{Weq})$ on category $\mathcal{M}$ with zero object, an object $X$ is trivial if $0\longrightarrow X$ is a weak equivalence, or, equivalently, $X\longrightarrow 0$ is a weak equivalence. It is cofibrant if $0\longrightarrow X$ is a cofibration, and it is fibrant if $X\longrightarrow 0$ is a fibration. An object is trivially cofibrant (respectively, trivially fibrant) if it is both trivial and cofibrant (respectively, fibrant). For a model structure on category $\mathcal{M}$ with zero object, Quillen’s homotopy category ${\rm Ho}(\mathcal{M})$ is the localization $\mathcal{M}[\operatorname{Weq}^{-1}]$. If $\mathcal{M}$ is an additive category, then ${\rm Ho}(\mathcal{M})$ is a pretriangulated category in the sense of [5]. It is not necessarily a triangulated category, in general.
A model structure on an exact category is exact ([14, 3.1]), if cofibrations are exactly inflations with cofibrant cokernel, and fibrations are exactly deflations with fibrant kernel. In this case, trivial cofibrations are exactly inflations with trivially cofibrant cokernel, and trivial fibrations are exactly deflations with trivially fibrant kernel. If $\mathcal{A}$ is an abelian category, then an exact model structure on $\mathcal{A}$ is exactly an abelian model structure in [19]. An exact model structure is projective if each object is fibrant, or equivalently, each trivially cofibrant object is projective ([14, 4.5, 4.6]). An exact model structure is trivial if each object is a trivial object. The homotopy category of a model structure is zero if and only if it is a trivial model structure ([18, Theorem 1.2.10(iv)]). We are only interested in exact model structures which are not trivial.
@@
M. Hovey [19] has established a one-one correspondence between abelian model structures and the Hovey triples in an abelian category. This has been extended to weakly idempotent complete exact category in J. Gillespie [14, Theorem 3.3] (see also [29, Theorem 6.9]). Namely, there is a one-to-one correspondence between exact model structures and the Hovey triples in weakly idempotent complete exact category $\mathcal{A}$, given by
@@
Thus, we identify a Hovey triple with an exact model structure, in weakly idempotent complete exact category. A Hovey triple $(\mathcal{C},\mathcal{F},\mathcal{W})$ is hereditary if $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})$ and  $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})$   are hereditary cotorsion pairs. An advantage of a hereditary Hovey triple is that the homotopy category of the corresponding exact model structure is just the stable category of a Frobenius category, and hence triangulated. See Lemma 1.3 below. Gillespie [15, Theorem 1.1] has constructed all the hereditary Hovey triples in an abelian category via two compatible complete cotorsion pairs.
@@ 1.4. Homotopy categories of exact model structures
$(1)$   ([29, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
@@
For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([29, 6.15]; compare [18, 6.2], or, cofibration sequences and fibration sequences in [24]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5]. The following nice result can be explicitly founded in Št́ovíček [29, Theorem 6.21]. See also Gillespie [14, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
@@
By Corollary 1.2, one only needs to show that the two triangulations in the same homotopy category also coincide. This is true, since by Theorem 1.3, the triangulation is determined by the cofiber sequences, which are by definition uniquely determined by conflations in $\mathcal{A}$ and by $\mathcal{W}$. See [29, Definition 6.15]. ∎
@@ 2. Abelian model structures arising from $\mathcal{PGF}_{n}$
([28, Theorem 4.9])  Let $R$ be a ring. Then $\mathcal{PGF}^{\perp}$ is thick; and $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair with kernel $\mathcal{P}$, and cogenerated by a set.
@@
([11, Theorem 7.4.6])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete and hereditary cotorsion pair, and cogenerated by a set.
The PGF version of Auslander - Buchweitz Theorem ([2, Theorem 1.1]; see also [17, Theorem 2.10]) can be found in G. Dalezios and I. Emmanouil [8].
@@
The following result describes the modules of finite PFG dimension $\leq n$. It is a theorem of Šaroch - Št́ovíček type: [28, Theorem 4.11] is the similar result, for Gorenstein flat modules.
@@
When $n=0$, the following result coincides with [28, Theorem 4.9].
@@
To prove that it is complete, it suffices to show that it is cogenerated by a set. By [28, Theorem 4.9] (cf. Lemma 2.1), the cotorsion pair  $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is cogenerated by a set, say $T$. Also the cotorsion pair  $(\mathcal{P}_{n},\mathcal{P}_{n}^{\perp})$ is cogenerated by a set, say $B$. Then $\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}=B^{\perp}\cap T^{\perp}=(B\cup T)^{\perp}$, i.e., the cotorsion pair  $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is cogenerated by the set $B\cup T$.
@@ 3. Abelian model structures arising from $\mathcal{GP}_{n}$
Auslander - Buchweitz Theorem is fundamental in Gorenstein homological algebra. See [2, Theorem 1.1], or more explicitly in [17, Theorem 2.10].
@@
The following is another theorem of Šaroch - Št́ovíček type (compare Theorem 2.5 and [28, Theorem 4.11]).
@@
It is clear that $\mathcal{P}_{n}\subseteq\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}$. Conversely, if $M\in\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}$, then by Lemma 3.2 there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. Since $K\in\mathcal{P}_{n-1}\subseteq\mathcal{GP}^{\perp}$ and $M\in\mathcal{GP}^{\perp}$, $G\in\mathcal{GP}^{\perp}$, and hence $G\in\mathcal{GP}\cap\mathcal{GP}^{\perp}=\mathcal{P}_{n}.$ Thus $M\in\mathcal{P}_{n}$.
@@
$(4)\Longrightarrow(1)$: Assume that  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ is an exact sequence with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{GP}$. Since $L\in\mathcal{P}_{n}\subseteq\mathcal{GP}_{n}$ and $N\in\mathcal{GP}\subseteq\mathcal{GP}_{n}$, it follows from [17, Theorem 2.20] that $M\in\mathcal{GP}_{n}$.
@@
Hence $\mathcal{B}=\mathcal{A}^{\perp}$ and $(\mathcal{A},\mathcal{B})$ is a cotorsion pair. It is hereditary, since $\mathcal{GP}_{n}$ is closed under taking kernels of epimorphisms (cf. [17, Theorem 2.20]); or equivalently $\mathcal{P}_{n}^{\perp}$, and hence $\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$, is closed under taking cokernels of monomorphisms.
@@
We do not need to consider the corresponding results for $\mathcal{GP}_{n}$ for Gorenstein ring $R$, by using the cortorsion pair  $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, since in this case one has $\mathcal{GP}=\mathcal{PGF}$ (cf. [20, Proposition 9]).

### v4

@@ Abstract.
Keywords: projectively coresolved Gorenstein flat (PGF) module; module of finite PGF (Gorenstein projective, Gorenstein flat) dimension; (complete, hereditary) cotorsion pair; abelian (exact) model structure; weakly idempotent complete exact category; finitistic dimension
@@ Introduction
This kind of PGF modules enjoys pleasant properties. It is not clear whether a Gorenstein projective module is Gorenstein flat; but PGF modules provide a common refinement of Gorenstein projective modules and Gorenstein flat modules: by definition they are Gorenstein flat, and they also prove to be Gorenstein projective ([32, Theorem 4.4]). The full subcategory $\mathcal{PGF}$ of $R$-Mod consisting of the PGF modules is a Frobenius category with $\mathcal{P}$, the full subcategory of projective $R$-modules, as the class of projective-injective objects, and the stable triangulated category $\mathcal{PGF}/\mathcal{P}$ is the homotopy category of infinite different abelian or exact model structures, in the sense of D. Quillen [28], M. Hovey [21], and J. Gillespie [15].
What interesting and important are, all the three classes $\mathcal{PGF}$, $\mathcal{GP}$, and $\mathcal{GF}$, are the left parts of cotorsion pairs in $R$-Mod, where $\mathcal{GP}$ (respectively, $\mathcal{GF}$) is the full subcategory of $R$-Mod consisting of the Gorenstein projective (respectively, Gorenstein flat) modules. A. Beligiannis and I. Reiten [5, X, Theorem 2.4] prove that $(\mathcal{GP},\ \mathcal{GP}^{\perp})$ is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra. Šaroch and Št́ovíček [32, Theorem 4.9, Corollary 4.12] prove that $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ and $(\mathcal{GF},\ \mathcal{EC}\cap\mathcal{PGF}^{\perp})$ are complete and hereditary cotorsion pairs in $R$-Mod, for any ring $R$. By a result of A. Iacob [22, Proposition 9] (see also [25, Proposition 4.11]), if $R$ is a Gorenstein ring (i.e., $R$ is a left and right noetherian ring and the injective dimensions of ${}_{R}R$ and $R_{R}$ are finite), then $\mathcal{PGF}=\mathcal{GP}$. Thus, in this case, the cotorsion pair $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ reads as $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, which has been given by Hovey [21, Theorem 8.3]. By the Hovey correspondence, all these three kinds of cotorsion pairs induce abelian model structures in $R$-Mod.
@@
are complete and hereditary cotorsion pairs in $R$-Mod (see [12, Theorem 7.4.6], and [26, Theorem 3.4(2)], respectively), where $\mathcal{P}_{n}$ (respectively, $\mathcal{F}_{n})$ is the class of $R$-modules of projective (respectively, flat) dimension $\leq n$.
@@
is a complete and hereditary cotorsion pair in $R$-Mod, if $R$ is an Artin algebra, where $\mathcal{GP}_{n}$ is the class of $R$-modules of Gorenstein projective dimension $\leq n$ (see Theorem 3.4). R. El Maaouy [25, Theorem A] has been proved that there is a complete and hereditary cotorsion pair in $R$-Mod
@@
Enochs, Jenda [12, Theorem 7.4.6]
@@
Enochs, Jenda [12, Proposition 7.4.3]
@@ 1.1. Modules of finite Gorenstein dimension
Following [32], an $R$-module $M$ is projectively coresolved Gorenstein flat (PGF for short), if it is a syzygy of an exact sequence of projective modules which remains exact after tensoring with any right injective $R$-module. Denote by $\mathcal{PGF}$ the subcategory of PGF modules. By definition $\mathcal{PGF}\subseteq\mathcal{GF}.$ It is proved in [32, Theorem 4.4] that $\mathcal{PGF}\subseteq\mathcal{GP}.$ However, it is not clear whether a Gorenstein projective module is PGF. In fact, $\mathcal{PGF}=\mathcal{GP}$ if and only if $\mathcal{GP}\subseteq\mathcal{GF}$. Denoted by ${\rm PGFd}M$ the PGF dimension of module $M$ (see [8]). Denote by $\mathcal{PGF}^{<\infty}$ the class of modules of finite PGF dimension. For a non-negative integer $n$, let $\mathcal{PGF}_{n}$ be the subcategory of $R$-modules $M$ with ${\rm PGFd}M\leq n$. Thus one has
@@ 1.2. Cotorsion pairs in exact categories
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is cogenerated by a set if $\mathcal{Y}=\mathcal{S}^{\perp}$, where $\mathcal{S}$ is a set of modules (or, equivalently, cogenerated by a single module). Any cotorsion pair in $R$-Mod cogenerated by a set is complete ([13, Theorem 10]; also [18, Theorem 6.11(b)]: but in [18] “cogenerated” is called “generated”). This result has been generalized to Grothendieck category with enough projective objects in [21, Theorem 2.4].
A cotorsion pair $(\mathcal{C},\mathcal{F})$ is hereditary, if $\mathcal{C}$ is closed under the kernel of deflations, and $\mathcal{F}$ is closed under the cokernel of inflations. For a cotorsion pair $(\mathcal{C},\mathcal{F})$ in an abelian category $\mathcal{A}$ with enough projective objects and injective objects, these two conditions of the heredity are equivalent; and they are equivalent to ${{\rm Ext}}^{2}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$ and $F\in\mathcal{F};$ and also to ${{\rm Ext}}^{i}_{\mathcal{A}}(\mathcal{C},\mathcal{F})=0$ for $C\in\mathcal{C}$, $F\in\mathcal{F}$, and $i\geq 2$. See [14, Proposition 1.2.10]. This result holds also for a complete cotorsion pair in a weakly idempotent complete exact category. See [33, Lemma 6.17] (also [31]).
@@ 1.3. Model structures
For a model structure $(\operatorname{CoFib}$,  $\operatorname{Fib}$,  $\operatorname{Weq})$ on category $\mathcal{M}$ with zero object, an object $X$ is trivial if $0\longrightarrow X$ is a weak equivalence, or, equivalently, $X\longrightarrow 0$ is a weak equivalence. It is cofibrant if $0\longrightarrow X$ is a cofibration, and it is fibrant if $X\longrightarrow 0$ is a fibration. An object is trivially cofibrant (respectively, trivially fibrant) if it is both trivial and cofibrant (respectively, fibrant). For a model structure on category $\mathcal{M}$ with zero object, Quillen’s homotopy category ${\rm Ho}(\mathcal{M})$ is the localization $\mathcal{M}[\operatorname{Weq}^{-1}]$. If $\mathcal{M}$ is an additive category, then ${\rm Ho}(\mathcal{M})$ is a pretriangulated category in the sense of [5]. It is not necessarily a triangulated category, in general. A model structure is trivial if each object is a trivial object. The homotopy category of a model structure is zero if and only if it is a trivial model structure ([20, Theorem 1.2.10(iv)]). We are only interested in model structures which are not trivial.
A model structure on an exact category is exact ([15, 3.1]), if cofibrations are exactly inflations with cofibrant cokernel, and fibrations are exactly deflations with fibrant kernel. In this case, trivial cofibrations are exactly inflations with trivially cofibrant cokernel, and trivial fibrations are exactly deflations with trivially fibrant kernel. If $\mathcal{A}$ is an abelian category, then an exact model structure on $\mathcal{A}$ is precisely an abelian model structure in [21]. An exact model structure is projective if each object is fibrant, or equivalently, each trivially cofibrant object is projective ([15, 4.5, 4.6]).
@@
M. Hovey [21] has established a one-one correspondence between abelian model structures and the Hovey triples in an abelian category. This has been extended to weakly idempotent complete exact category in J. Gillespie [15, Theorem 3.3] (see also [33, Theorem 6.9]). Namely, there is a one-to-one correspondence between exact model structures and the Hovey triples in weakly idempotent complete exact category $\mathcal{A}$, given by
@@
Thus, we identify a Hovey triple with an exact model structure, in weakly idempotent complete exact category. A Hovey triple $(\mathcal{C},\mathcal{F},\mathcal{W})$ is hereditary if $(\mathcal{C}\cap\mathcal{W},\ \mathcal{F})$ and  $(\mathcal{C},\ \mathcal{F}\cap\mathcal{W})$   are hereditary cotorsion pairs. An advantage of a hereditary Hovey triple is that the homotopy category of the corresponding exact model structure is just the stable category of a Frobenius category, and hence triangulated. See Theorem 1.3 below. Gillespie [16, Theorem 1.1] has constructed all the hereditary Hovey triples in an abelian category via two compatible hereditary complete cotorsion pairs.
Furthermore, H. Nakaoka and Y. Palu [NP] introduce an extriangulated category, which is a common generalization of an exact category and a triangulated category, and generalize the Hovey correspondence to this setting up; moreover, remarkably, the homotopy category is proved to be triangulated, even if the two compatible complete cotorsion pairs are not necessarily hereditary. See [27, Theorem 6.20], and [17, Theorem 6.34].
@@ 1.4. Homotopy categories of exact model structures
$(1)$   ([33, 6.9(b)])  $\operatorname{Weq}=\{d\circ i\ |\ i\ \mbox{is an inflation,}\ {\rm Coker}i\in\mathcal{W},\ d\ \mbox{is a deflation,}\ {\rm Ker}d\in\mathcal{W}\}.$
@@
For an exact model structure on weakly idempotent complete exact category $\mathcal{A}$, with cofiber sequences and fiber sequences in ${\rm Ho}(\mathcal{A})$ ([33, 6.15]; compare [20, 6.2], or, cofibration sequences and fibration sequences in [28]), the homotopy category ${\rm Ho}(\mathcal{A})$ is a pretriangulated category in the sense of [5]. The following nice result can be explicitly founded in Št́ovíček [33, Theorem 6.21]. See also Gillespie [15, Proposition 4.4, Proposition 5.2] and Becker [4, Proposition 1.1.14].
@@
By Corollary 1.2, one only needs to show that the two triangulations in the same homotopy category also coincide. This is true, since by Theorem 1.3, the triangulation is determined by the cofiber sequences, which are by definition uniquely determined by conflations in $\mathcal{A}$ and by $\mathcal{W}$. See [33, Definition 6.15]. ∎
@@ 2. Abelian model structures arising from $\mathcal{PGF}_{n}$
([32, Theorem 4.9])  Let $R$ be a ring. Then $\mathcal{PGF}^{\perp}$ is thick; and $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is a complete and hereditary cotorsion pair with kernel $\mathcal{P}$, and cogenerated by a set.
@@
([12, Theorem 7.4.6])  Let $R$ be a ring, and $n$ a non-negative integer. Then $(\mathcal{P}_{n},\ \mathcal{P}_{n}^{\perp})$ is a complete and hereditary cotorsion pair, and cogenerated by a set.
The PGF version of Auslander - Buchweitz Theorem ([2, Theorem 1.1]; see also [19, Theorem 2.10]) can be found in G. Dalezios and I. Emmanouil [8].
@@
The following result describes the modules of finite PGF dimension $\leq n$. It is a theorem of Šaroch - Št́ovíček type: [32, Theorem 4.11] is the similar result, for Gorenstein flat modules.
@@
When $n=0$, the following result coincides with [32, Theorem 4.9].
@@
To prove that it is complete, it suffices to show that it is cogenerated by a set. By [32, Theorem 4.9] (cf. Lemma 2.1), the cotorsion pair  $(\mathcal{PGF},\ \mathcal{PGF}^{\perp})$ is cogenerated by a set, say $T$. Also the cotorsion pair  $(\mathcal{P}_{n},\mathcal{P}_{n}^{\perp})$ is cogenerated by a set, say $B$. Then $\mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp}=B^{\perp}\cap T^{\perp}=(B\cup T)^{\perp}$, i.e., the cotorsion pair  $(\mathcal{PGF}_{n},\ \mathcal{P}_{n}^{\perp}\cap\mathcal{PGF}^{\perp})$ is cogenerated by the set $B\cup T$.
@@ 3. Abelian model structures arising from $\mathcal{GP}_{n}$
Auslander - Buchweitz Theorem is fundamental in Gorenstein homological algebra. See [2, Theorem 1.1], or more explicitly in [19, Theorem 2.10].
@@
The following is another theorem of Šaroch - Št́ovíček type (compare Theorem 2.5 and [32, Theorem 4.11]).
@@
It is clear that $\mathcal{P}_{n}\subseteq\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}$. Conversely, if $M\in\mathcal{GP}^{\perp}\cap\mathcal{GP}_{n}$, then by Lemma 3.2 there is an exact sequence $0\longrightarrow K\longrightarrow G\longrightarrow M\longrightarrow 0$ with $G\in\mathcal{GP}$ and $K\in\mathcal{P}_{n-1}$. Since $K\in\mathcal{P}_{n-1}\subseteq\mathcal{GP}^{\perp}$ and $M\in\mathcal{GP}^{\perp}$, $G\in\mathcal{GP}^{\perp}$, and hence $G\in\mathcal{GP}\cap\mathcal{GP}^{\perp}=\mathcal{P}.$ Thus $M\in\mathcal{P}_{n}$.
@@
$(4)\Longrightarrow(1)$: Assume that  $0\longrightarrow M\longrightarrow L\longrightarrow N\longrightarrow 0$ is an exact sequence with $L\in\mathcal{P}_{n}$ and $N\in\mathcal{GP}$. Since $L\in\mathcal{P}_{n}\subseteq\mathcal{GP}_{n}$ and $N\in\mathcal{GP}\subseteq\mathcal{GP}_{n}$, it follows from [19, Theorem 2.20] that $M\in\mathcal{GP}_{n}$.
@@
Hence $\mathcal{B}=\mathcal{A}^{\perp}$ and $(\mathcal{A},\mathcal{B})$ is a cotorsion pair. It is hereditary, since $\mathcal{GP}_{n}$ is closed under taking kernels of epimorphisms (cf. [19, Theorem 2.20]); or equivalently $\mathcal{P}_{n}^{\perp}$, and hence $\mathcal{P}_{n}^{\perp}\cap\mathcal{GP}^{\perp}$, is closed under taking cokernels of monomorphisms.
@@
We do not need to consider the corresponding results for $\mathcal{GP}_{n}$ for Gorenstein ring $R$, by using the cortorsion pair  $(\mathcal{GP},\ \mathcal{P}^{<\infty})$, since in this case one has $\mathcal{GP}=\mathcal{PGF}$ (cf. [22, Proposition 9]).